The function f(x,y,z) = e^(xyz) is integrable over [0,2] × [0,2] × [0,2] .
1
Expert's answer
2020-03-21T17:03:48-0400

We first calculate the integral

"\int\limits_{0}^{2}e^{xyz} dz=\\
=\left\{\begin{matrix}
 \frac{e^{xyz}}{xy}|_0^2, for xy\neq 0 \\
 z|_0^2, otherwise
\end{matrix}\right.=\\
\left\{\begin{matrix}
 \frac{e^{2xy}}{xy}-\frac{1}{xy}, for xy\neq 0 \\
 2, otherwise
\end{matrix}\right. .\\"

Then find

"\int\limits_{0}^{2}
\left\{\begin{matrix}
 \frac{e^{2xy}}{xy}-\frac{1}{xy}, for xy\neq 0 \\
 2, otherwise
\end{matrix}\right. dy=\\

=\left\{\begin{matrix}
 \int\limits_{0}^{2}\frac{e^{2xy}-1}{xy}dy, for xy\neq 0 \\
 \int\limits_{0}^{2} 2dy, otherwise
\end{matrix}\right. =\\
=\left\{\begin{matrix}
 \int\limits_{0}^{2}\frac{e^{2xy}-1}{xy}dy, for xy\neq 0 \\
2y|^2_0, otherwise
\end{matrix}\right. =\\
=\left\{\begin{matrix}
 \int\limits_{0}^{2}\frac{e^{2xy}-1}{xy}dy, for xy\neq 0 \\
4, otherwise
\end{matrix}\right. ."

Then find

"\int\limits_{0}^{2}\left\{\begin{matrix}
 \int\limits_{0}^{2}\frac{e^{2xy}-1}{xy}dy, for xy\neq 0 \\
4, otherwise
\end{matrix}\right. dx=\\
=\left\{\begin{matrix}
 \int\limits_{0}^{2} \int\limits_{0}^{2}\frac{e^{2xy}-1}{xy}dxdy, for xy\neq 0 \\
\int\limits_{0}^{2} 4 dx, otherwise
\end{matrix}\right. =\\
=\left\{\begin{matrix}
 \int\limits_{0}^{2} \int\limits_{0}^{2}\frac{e^{2xy}-1}{xy}dxdy, for xy\neq 0 \\
4 x|_0^2, otherwise
\end{matrix}\right. =\\
=\left\{\begin{matrix}
 \int\limits_{0}^{2} \int\limits_{0}^{2}\frac{e^{2xy}-1}{xy}dxdy, for xy\neq 0 \\
8, otherwise
\end{matrix}\right. ."

Consider the function "\frac{e^{2xy}-1}{xy}" and find limit

"\lim\limits_{xy\mapsto 0} \frac{e^{2xy}-1}{xy}=2"

because

"\lim\limits_{t\mapsto 0} \frac{e^{t}-1}{t}=1" .

This means that the function is integrable and the integral is convergent.